# Improving Signed Propagation of Graph Neural Network Under Multiple Classes

### v6

@@ IV-B Using Signed Messages on Binary Classes
Figure 3: Node classification accuracy on six benchmark datasets. Vanilla GCN utilizes the original graph, while the coefficient of heterophilous edges is changed to -1 in signed GCN and to 0 in zero-weight GCN, respectively. Here, signed GCN (+ calib) further employs two types of calibration, which is introduced in Section V
IV-C Empirical Analysis
@@
Through Corollary 4.2 and Eq. IV-B, we prove that signed propagation is generally advantageous when it is applied to binary class graphs. Now, we examine whether this finding is valid on multi-class graphs with more than two classes. In Figure 3, we measure the node classification accuracy of GCN [2] on nine benchmark graphs (the statistical details of the datasets are shown in Table II). Starting from the original GCN called vanilla GCN, we construct two variants; one variant replaces heterophilic edges with -1 (called signed GCN), and the other assigns weight 0 to heterophilic edges (called zero-weight GCN). Since we identified all heterophilous edges, the error ratio is zero. In this case, $Z$ $(=1-b_{i})$ derived in Corollary 4.2 has a non-negative value regardless of the local homophily ratio $b_{i}$, meaning that signed GCN outperforms zero-weight GCN. However, Figure 3 shows that zero-weight GCN generally outperforms signed GCN ($Z\leq 0$). To investigate this phenomenon, we extend the theorems to a multi-class scenario and introduce two types of calibration below.
IV-D Signed Messaging on Multi-Class Graphs
@@
$\begin{gathered}\mathbb{E}(h^{(0)}_{i}|y_{i})=(\mu,\phi,\theta),\end{gathered}$ (14)
where $\mu$ represents the scale of a vector. The direction is determined by two angles $\phi$ and $\theta$. As shown in Figure 1b, this equation satisfies the origin symmetry as in the binary case since $(\mu,\pi/2,0)=(-\mu,\pi/2,0)$. Based on this, we introduce a novel understanding by extending Eq. IV-B and Eq. IV-B as follows:
Theorem 4.3 (Signed GCN).
Let us assume $y_{i}=0$. For simplicity, we fix the coordinates of the ego $(\mu,\pi/2,0)$ as $k$, and denote the aggregated neighbors as $k^{\prime}$ = $(\mu^{\prime},\phi^{\prime},\theta^{\prime})$, where $0\leq\mu^{\prime}\leq\mu$, $\phi^{\prime}=\pi/2$, and $0\leq\theta^{\prime}\leq 2\pi$ (please see Figure 1c). Then, the expectation of $h^{(1)}_{i}$ is defined to be:
$\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={k\over d_{i}+1}+$
$\sum_{j\in\mathcal{N}_{i}}\left({k(1-e)-ke\over d_{ij}}b_{i}+{-k^{\prime}(1-e)+k^{\prime}e\over d_{ij}}(1-b_{i})\right)$
$={(1-2e)\{b_{i}k+(b_{i}-1)k^{\prime}\}d^{\prime}_{i}+k\over d_{i}+1}$ (15)
Assuming the scale of all vectors as $\mu$, the aggregated vector $k^{\prime}$ always satisfies $|k^{\prime}|\leq\mu$ considering the degree-normalization and homophily coefficient ($1-b_{i}\leq 1$).
Theorem 4.4 (Zero-weight GCN).
@@
$\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={k\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}\left({k(1-e)\over d_{ij}}b_{i}+{k^{\prime}e\over d_{ij}}(1-b_{i})\right)$
=$={\{(1-e)b_{i}k+e(1-b_{i})k^{\prime}\}d^{\prime}_{i}+k\over d_{i}+1}$ (16)
Similar to the Corollary 4.2, we can compare the separability of the two methods based on their coefficients.
@@
The difference of separation powers ($Z$) between signed GCN and zero-weight GCN in multi-class graphs is given by:
@@
$=-eb_{i}k+(1-e)(b_{i}-1)k^{\prime}$ (17)
Based on the above equation, we can induce the following conditional statement:
$\begin{gathered}Z\in\begin{cases}(1-e-b_{i})k,&\text{if}\,\,\,k^{\prime}=-k\\ \{-2eb_{i}-(1-e-b_{i})\}k,&\text{if}\,\,\,k^{\prime}=k\end{cases}\end{gathered}$ (18)
Specifically, as the distribution of aggregated neighbors $k^{\prime}$ moves away from $k$ (e.g., $k^{\prime}=-k$), $Z=1-e-b_{i}$ becomes identical to Eq. 10. This leads to the following equation,
@@
indicating that signed propagation outperforms the zero-weighting scheme. However, as $k^{\prime}$ gets closer to $k$, we claim that $Z=-2eb_{i}+e+b_{i}-1$ tends to be negative, where signed propagation may achieve dismal performance as below:
@@
$=\left[{-eb_{i}^{2}-e^{2}b_{i}+e^{2}+b_{i}^{2}\over 2}-1\right]^{1}_{e,b_{i}=0}=-1$ (20)
Intuitively, the probability of being $\cos(k^{\prime},k)=-1$ is inversely proportional to the number of entire classes. Thus, we can infer that zero-weight GCN generally outperforms signed one given multi-class datasets. As shown in Fig. 3, signed GCN outperforms zero-weight GCN only on the Pubmed dataset which has the smallest number of classes (Table 3) among the datasets. Thus, we propose to manually adjust the messages from the neighbors ($k^{\prime}$) based on their similarity, denoted as edge weight calibration. The details are introduced in Section V-A, which guides $k^{\prime}$ to get closer to $-k$ as in the upper case of Eq. 18 ($Z=1-e-b_{i}$).
Parameter update (Signed propagation contributes to the ego-neighbor separation, but it also increases the prediction uncertainty under multiple class environments) Uncertainty management, which is closely related to the entropy [33], is vital in securing confident predictions [34, 35]. Here, we focus on the conflicting evidence [36, 37] which ramps up the entropy of outputs. One can easily measure the uncertainty of a prediction ($\widehat{y}_{i}$) using Shannon’s entropy [38] as below:
$H(\widehat{y}_{i})=-\sum^{C}_{j=1}\widehat{y}_{i,j}log_{c}\widehat{y}_{i,j}$ (21)
@@
Figure 4: Visualization of parameter update using the Dirichlet distribution (central side means higher prediction uncertainty). (a) Binary class case and (b) Multi-class case. In both cases, signed messages separate the ego and neighbors. However, in (b) multi-class case, the uncertainty of neighbors $j$ and $k$ increases
Theorem 4.6 (Uncertainty).
Signed messaging contributes to ego-neighbor separation (Eq. IV-B). However, as the training epoch $t$ proceeds, the expectation of signed GNN prediction ($\widehat{y}_{s}$) exhibits higher uncertainty $H(\mathbb{E}[\widehat{y}_{s}])$ compared to that ($\widehat{y}_{p}$) of plain GNN $H(\mathbb{E}[\widehat{y}_{p}])$ as below:
$\begin{gathered}H(\mathbb{E}[\widehat{y}^{(t+1)}_{s}])-H(\mathbb{E}[\widehat{y}^{(t+1)}_{p}])>H(\mathbb{E}[\widehat{y}^{t}_{s}])-H(\mathbb{E}[\widehat{y}^{t}_{p}])\end{gathered}$ (22)
Proof. Contrary to Eq. IV-B, we can infer that the true class probability ($k$) of node $p$ ($\widehat{y}_{p,k}$) increases, while other class probabilities $\widehat{y}_{p,o}$ ($o\neq k$) decrease for plane GNN as follows:
$\begin{gathered}\widehat{y}^{(t+1)}_{p}\in\begin{cases}\widehat{y}^{t}_{p,k}-\eta\nabla_{p}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{k}>\widehat{y}^{t}_{p,k}\\ \widehat{y}^{t}_{p,o}-\eta\nabla_{p}\mathcal{L}_{nll}(Y_{i},\widehat{Y}_{i})_{o}<\widehat{y}^{t}_{p,o}\,\,\,\,\,\forall\,\,o\neq k\end{cases}\end{gathered}$ (23)
As shown in Eq. IV-B, the gradient of node $s$ will have a reversed sign compared to node $p$. Thus, as the training epoch increases, $\widehat{y}_{p,k}$ will converge to 1 resulting in the decrease of $H(\mathbb{E}[\widehat{y}_{p}])$. Conversely, $\widehat{y}_{s,k}$ gets closer to 0, which may fail to generate a highly confident prediction and lead to a surge of uncertainty. Thus, one can infer that $H(\mathbb{E}[\widehat{y}^{(t+1)}_{s}])-H(\mathbb{E}[\widehat{y}^{(t+1)}_{p}])>H(\mathbb{E}[\widehat{y}^{t}_{s}])-H(\mathbb{E}[\widehat{y}^{t}_{p}])$ as $t\rightarrow\infty$, which completes the proof. ∎
In Figure 4, we visualize the impact of signed propagation using the Dirichlet distribution. In both figures, signed messages contribute to the separation. However, in Fig 4(b), the simultaneous increase in other class probabilities leads to the increment of conflict evidence. The solution to this will be introduced in §$\S$ V-B as a confidence calibration.
V Methodology
We introduced the limitations of signed propagation in the multi-class scenario from two perspectives; message-passing and parameter update. Here, we propose two types of calibration, both of which belong to the self-training method [15, 39].
V-A (Message-Passing) Edge Weight Calibration
@@ V Methodology
$\forall(i,j)\in A,\,\,\tilde{A}_{ij}={\cos(h^{l}_{i},h^{l}_{j})+1\over 2}\cdot A_{ij}$ (24)
One important thing is that we should also consider the sign of a message from multi-hop away nodes, which may be inconsistent through which information is transmitted. For example, we assume three nodes $i,j,k$ that are connected serially and their edge weights are $s_{ij}$ and $s_{jk}$, respectively. If $y_{i}=y_{j}$, then $s_{ij}=1$, otherwise $s_{ij}=-1$. Under binary class, multi-hop edge weight $s_{ij}\cdot s_{jk}$ is always consistent with the node label. However, under multiple classes, if the labels of three nodes are $y_{i}=0$, $y_{j}=1$, and $y_{k}=2$, $k$ is trained to be similar to $i$ even though $i$ and $k$ are in different classes since $s_{ij}\cdot s_{jk}=-1\times-1=1$. To address this, we modify the above equation using Jumping Knowledge [40], assuming $l$-hop message-passing as shown below:
$\forall(i,j)\in\sum^{L}_{l=1}A^{L},\,\,\tilde{A}_{ij}={\cos(h^{l}_{i},h^{l}_{j})+1\over 2}\cdot A_{ij}$ (25)
@@
$\begin{gathered}\mathbb{E}\left[|k^{\prime}_{edge}|\right]={|k|\over d}\sqrt{{c\over 4}+\sum^{c}_{j=0}\cos({2\pi j\over c})^{2}}>\mathbb{E}\left[|k^{\prime}|\right]\end{gathered}$ (26)
In Eq. 4.5, we can substitute $k^{\prime}_{edge}$ into $k$ and derive $\mathbb{E}\left[Z\right]\leq\mathbb{E}\left[Z_{edge}\right]$ which demonstrates an increase in separability.
Proof. Similar to Eq. 4.3, we assume that $|k^{\prime}|\leq\mu$ and $|k^{\prime}_{edge}|\leq\mu$. Since the discrimination power is determined based on the distance from the decision boundary, we can substitute $k^{\prime}$ and $k^{\prime}_{edge}$ in terms of $k$. Specifically, Eq. 4.5 can be redefined as below:
$Z$ $=-eb_{i}k+(1-e)(b_{i}-1)\underline{\cos(k^{\prime},k)|k^{\prime}|}$
$Z$ =${}_{edge}=-eb_{i}k+(1-e)(b_{i}-1)\underline{\cos(k^{\prime}_{edge},k)|k^{\prime}_{edge}|}$ (27)
Now, let’s assume the number of classes is $c$, the average degree of nodes is $d$, and every class has the same population. Based on the above equation, one can compare the scale and direction between $\cos(k^{\prime},k)|k|$ and $\cos(k^{\prime}_{edge},k)|k^{\prime}_{edge}|$ since other variables are the same. As $k$ is a fixed vector (ego), the expectation of $\cos(k^{\prime},k)|k^{\prime}|$ can be represented as,
${|k|\over d}\sqrt{\Big{\{}\sum^{c}_{j=0}\cos({2\pi j\over c})\Big{\}}^{2}+\Big{\{}\sum^{c}_{j=0}\sin({2\pi j\over c})\Big{\}}^{2}}$ (28)
Under an identical class distribution, we can imagine an n-sided polygon inscribed in a circle, and the sum of the elements of the neighbor vectors $k^{\prime}$ is 0. By replacing $c^{\prime}=2\pi/c$, we can induce $\cos(k^{\prime}_{edge},k)|k^{\prime}_{edge}|$ as follows:
$\mathbb{E}\left[\cos(k^{\prime}_{edge},k)|k^{\prime}_{edge}|\right]$
$={|k|\over d}\sqrt{\Big{\{}\sum^{c}_{j=0}-\cos(jc^{\prime})^{2}\Big{\}}^{2}+\Big{\{}\sum^{c}_{j=0}-\cos(jc^{\prime})\sin(jc^{\prime})\Big{\}}^{2}}$
$={|k|\over d}\sqrt{\Big{\{}\sum^{c}_{j=0}{-1-\cos(2\pi jc^{\prime})\over 2}\Big{\}}^{2}+\Big{\{}\sum^{c}_{j=0}-{\sin(2\pi jc^{\prime})\over 2}\Big{\}}^{2}}$
$={|k|\over d}\sqrt{{c\over 4}+\sum^{c}_{j=0}\cos(jc^{\prime})^{2}}>\mathbb{E}\left[\cos(k^{\prime},k)|k^{\prime}|\right]$ (29)
Given that $\mathbb{E}\left[\cos(k^{\prime}_{edge},k)|k^{\prime}_{edge}|\right]$ has some positive value $\alpha$, we can infer the following inequality:
$\mathbb{E}\left[Z_{edge}\right]=\mathbb{E}\left[Z\right]+\alpha>\mathbb{E}\left[Z\right]\,\,\,\,\,\,\,\,\,\,\,\,(\alpha>0)$ (30)
which completes the proof. ∎
@@
Remark. The edge weight calibration also enhances the quality of positive GNNs like GCN [2] and GAT [3] by reflecting the cosine similarity of two nodes [6] as below.
V-A1 Edge weight calibration for GCN
@@
We first explain how edge weight calibration improves the quality of positive GNNs. For brevity, we take the aggregation scheme of GCN [2]. Then, the expectation of node $i$ after message-passing $h^{(1)}_{i}$ becomes,
$\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={k\over d_{i}+1}$
$+\sum_{j\in\mathcal{N}_{i}}\left({k(1-e)+ke\over d_{ij}}b_{i}+{k^{\prime}(1-e)+k^{\prime}e\over d_{ij}}(1-b_{i})\right)$
$={k\over d_{i}+1}+{\{kb_{i}+k^{\prime}(1-b_{i})\}d^{\prime}\over d_{i}+1}$ (31)

### v7

@@ 4.2. Using Signed Messages on Binary Classes
Figure 3. Node classification accuracy on six benchmark datasets. Vanilla GCN utilizes the original graph, while the coefficient of heterophilous edges is changed to -1 in signed GCN and to 0 in zero-weight GCN, respectively. Here, signed GCN (+ calib) further employs two types of calibration, which is introduced in Section 5
4.3. Empirical Analysis
@@
Through Corollary 4.1 and Eq. 13, we prove that signed propagation is generally advantageous when it is applied to binary class graphs. Now, we examine whether this finding is valid on multi-class graphs with more than two classes. In Figure 3, we measure the node classification accuracy of GCN (Kipf and Welling, 2016) on nine benchmark graphs (the statistical details of the datasets are shown in Table 2). Starting from the original GCN called vanilla GCN, we construct two variants; one variant replaces heterophilic edges with -1 (called signed GCN), and the other assigns weight 0 to heterophilic edges (called zero-weight GCN). Since we identified all heterophilous edges, the error ratio is zero. In this case, $Z$ $(=1-b_{i})$ derived in Corollary 4.1 has a non-negative value regardless of the local homophily ratio $b_{i}$, meaning that signed GCN outperforms zero-weight GCN. However, Figure 3 shows that zero-weight GCN generally outperforms signed GCN ($Z\leq 0$). To further investigate this phenomenon, we extend the theorems to the multi-class scenario and identify some drawbacks. We then introduce two types of calibration to enhance separability while reducing uncertainty.
4.4. Signed Messaging on Multiple Classes
@@
(14) $\begin{gathered}\mathbb{E}(h^{(0)}_{i}|y_{i})\sim(\mu,\phi,\theta)\end{gathered}$
where $\mu$ represents the scale of a vector. The direction is determined by two angles $\phi$ and $\theta$. As shown in Figure 1b, this equation satisfies the origin symmetry as in the binary case since $(\mu,\pi/2,0)=(-\mu,\pi/2,0)$. Based on this, we introduce a novel understanding by extending Eq. 7 and Eq. 8 as follows:
Lemma 4.2 (Signed GCN).
Let us assume $y_{i}=0$. For simplicity, we assume the coordinates of the ego $(\mu,\pi/2,0)$ as $k$, and denote the aggregated neighbors as $k^{\prime}$ = $(\mu^{\prime},\phi^{\prime},\theta^{\prime})$, where $0\leq\mu^{\prime}\leq\mu$, $\phi^{\prime}=\pi/2$, and $0\leq\theta^{\prime}\leq 2\pi$ as explained in Figure 1. Then, the expectation of $h^{(1)}_{i}$ is defined to be:
(15) $\mathbb{E}^{m}_{s}(h^{(1)}_{i}|v_{i},d_{i})={(1-2e)\{b_{i}k+(b_{i}-1)k^{\prime}\}d^{\prime}_{i}+k\over d_{i}+1}$
Given the scale of all vectors is $\mu$, the aggregated vector $k^{\prime}$ always satisfies $|k^{\prime}|\leq\mu$ considering the degree-normalization and homophily coefficient ($1-b_{i}\leq 1$).
Lemma 4.3 (Zero-weight GCN).
@@
(16) $\mathbb{E}^{m}_{z}(h^{(1)}_{i}|v_{i},d_{i})={\{(1-e)b_{i}k+e(1-b_{i})k^{\prime}\}d^{\prime}_{i}+k\over d_{i}+1}$
$\bullet$ Derivation can be found in Appendix B-1 $\sim$ B-2.
Similar to the Corollary 4.1, we can compare the separability of the two methods under multiple classes.
@@
The difference of separation powers $Z=\mathbb{E}^{m}_{s}(\cdot)-\mathbb{E}^{m}_{z}(\cdot)$ between signed GCN and zero-weight GCN in multi-class graphs is given by:
@@
(17) $=-eb_{i}k+(1-e)(b_{i}-1)k^{\prime}$
Accordingly, we can induce the following conditional statement:
(18) $\begin{gathered}Z\in\begin{cases}(1-e-b_{i})k,&\text{if}\,\,\,k^{\prime}=-k\\ \{-2eb_{i}-(1-e-b_{i})\}k,&\text{if}\,\,\,k^{\prime}=k\end{cases}\end{gathered}$
$\bullet$ More details are introduced in Appendix B-3.
Similar to the binary class, as the distribution of aggregated neighbors $k^{\prime}$ diverges from $k$ (e.g., $k^{\prime}=-k$), $Z$ $(=1-e-b_{i})$ becomes identical to Eq. 10, which converges to zero as follows:
@@
This means selecting a proper propagation scheme depends on $b_{i}$ and $e$. However, as $k^{\prime}$ gets closer to $k$, we can infer that $Z$ $(=-2eb_{i}+e+b_{i}-1)$ tends to be negative as below:
@@
where signed propagation may achieve dismal performance. Intuitively, the probability of being $\cos(k^{\prime},k)=-1$ is inversely proportional to the number of entire classes. Thus, we conclude that zero-weight GCN generally outperforms signed one under multi-class datasets. As shown in Fig. 3, signed GCN outperforms zero-weight GCN only on the Pubmed dataset which has the smallest number of classes (Table 3) among the datasets. Thus, we propose to manually adjust the messages from the neighbors ($k^{\prime}$) based on their similarity, denoted as edge weight calibration. The details are introduced in Section 5.1, which guides $k^{\prime}$ to get closer to $-k$ as in the upper case of Eq. 18 ($Z=1-e-b_{i}$).
Parameter update (Although signed propagation contributes to the ego-neighbor separation, it also increases the prediction uncertainty under multiple class environments) Uncertainty management, which is closely related to the entropy (Liu et al., 2022), is vital in securing confident predictions (Moon et al., 2020; Mukherjee and Awadallah, 2020). Here, we focus on the conflicting evidence (Perry, 2013; Zhao et al., 2020) which ramps up the entropy of outputs. One can easily measure the uncertainty of a prediction ($\widehat{y}_{i}$) using Shannon’s entropy (Shannon, 1948) as below:
(21) $H(\widehat{y}_{i})=-\sum^{C}_{j=1}\widehat{y}_{i,j}log_{c}\widehat{y}_{i,j}$
@@
Figure 4. Visualization of parameter update using the Dirichlet distribution (central side means higher prediction uncertainty). (a) Binary class case and (b) Multi-class case. In both cases, signed messages separate the ego and neighbors. However, in (b) multi-class case, the uncertainty of neighbors ($j$ and $k$) connected with signed edges increases
Corollary 4.5 (Prediction Uncertainty).
Signed messaging contributes to ego-neighbor separation (Eq. 13). Nonetheless, in the case of multiple classes, the expected prediction uncertainty when using the signed propagation, $H(\mathbb{E}[\widehat{y}_{s}])$, remains higher than that of positive/zero-weighted, $H(\mathbb{E}[\widehat{y}_{p}])$, as the training epoch ($t$) progresses:
(22) $\begin{gathered}\lim_{t\to\infty}\{H(\mathbb{E}[\widehat{y}^{t}_{s}])-H(\mathbb{E}[\widehat{y}^{t}_{p}])\}>0\end{gathered}$
$\bullet$ The proof is given in Appendix C.
For ease of understanding, in Figure 4, we visualize the impact of signed propagation using the Dirichlet distribution. In both figures, signed messages contribute to the separation. However, in Fig 4(b), the simultaneous increase in other class probabilities leads to the increment of conflict evidence. The solution to this will be introduced in §$\S$ 5.2 as a confidence calibration.
5. Methodology
We introduced the limitations of signed propagation in the multi-class scenario from two perspectives; message-passing and parameter update. Here, we propose two types of calibration, both of which belong to the self-training method (Guo et al., 2017; Yang et al., 2021), respectively.
5.1. (Message-Passing) Edge Weight Calibration
@@ 5. Methodology
(23) $\forall(i,j)\in A,\,\,\tilde{A}_{ij}={\cos(h^{l}_{i},h^{l}_{j})+1\over 2}\cdot A_{ij}$
We should also consider the sign of a message from multi-hop away nodes, which may be inconsistent through which information is transmitted. For example, we assume three nodes $i,j,k$ that are connected serially and their edge weights are $s_{ij}$ and $s_{jk}$, respectively. If $y_{i}=y_{j}$, then $s_{ij}=1$, otherwise $s_{ij}=-1$. Under binary class, multi-hop edge weight $s_{ij}\cdot s_{jk}$ is always consistent with the node label. However, under multiple classes, if the labels of three nodes are $y_{i}=0$, $y_{j}=1$, and $y_{k}=2$, $k$ is trained to be similar to $i$ even though $i$ and $k$ are in different classes since $s_{ij}\cdot s_{jk}=-1\times-1=1$. To address this, we modify the above equation using Jumping Knowledge (Xu et al., 2018b; Lei et al., 2022), assuming $l$-hop message-passing as shown below:
(24) $\forall(i,j)\in\sum^{L}_{l=1}A^{L},\,\,A_{ij}={\cos(h^{l}_{i},h^{l}_{j})+1\over 2}\cdot A_{ij}$
@@
(25) $\begin{gathered}\mathbb{E}\left[|k^{\prime}_{edge}|\right]={|k|\over d}\sqrt{{c\over 4}+\sum^{c}_{j=0}\cos({2\pi j\over c})^{2}}>\mathbb{E}\left[|k^{\prime}|\right]\end{gathered}$
where we prove that an increase in distance from the decision boundary $\mathbb{E}\left[|k^{\prime}_{edge}|\right]>\mathbb{E}\left[|k^{\prime}|\right]$ indicates improved separability.
$\bullet$ Proof is provided in Appendix D.
@@
Below, we demonstrate that edge weight calibration also improves the performance of positive GNNs, such as GCN (Kipf and Welling, 2016) and GAT (Velickovic et al., 2017), by reflecting the cosine similarity of two nodes (Brody et al., 2021).
5.1.1. Edge weight calibration for positive GNN
@@
We explain how edge weight calibration improves the quality of positive GNNs. For brevity, we take the aggregation scheme of GCN (Kipf and Welling, 2016). Then, the expectation of node $i$ after message-passing $h^{(1)}_{i}$ becomes,
(26) $\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})={k\over d_{i}+1}+{\{kb_{i}+k^{\prime}(1-b_{i})\}d^{\prime}\over d_{i}+1}$